# Knowledge-Driven Federated Graph Learning on Model Heterogeneity

### v3

@@ 3.4. Knowledge-Aware Model Aggregation
Given the heterogeneity among clients, the learning efficiency of each client’s model varies accordingly. Traditional FL methods typically employ an averaging aggregation strategy (McMahan et al., 2017), which does not take into account the model heterogeneity scenario. This oversight results in conflicted models aggregation and suboptimal overall performance. To address it, we propose a Knowledge-Aware Model Aggregation mechanism that dynamically assigns weights based on learned knowledge quality of each client’s model. Through dynamic weighting, well-informed clients is protected from the adverse effects of clients with limited knowledge. Simultaneously, it channels greater guidance from knowledgeable clients to less-informed ones, improving convergence and robustness.
@@ 5.1.1. Datasets
Experiments are conducted utilizing 8 commonly employed benchmark datasets in graph learning. These consist of three small-scale citation networks (Cora, CiteSeer, PubMed) (Yang et al., 2016), two medium-scale co-authorship networks (CS, Physics) (Shchur et al., 2018), two medium-scale user-item datasets (Photo, Computer) (McAuley et al., 2015), and one large-scale open graph benchmark dataset (ogbn-arxiv) (Hu et al., 2020). Further particulars regarding these datasets are delineated in Table 6.
@@ 5.1.2. Baselines
We compare FedGKC with heterogeneous federation strategies. Given the absence of strategies for MHtFGL, we re-implemented these methods under FGL conditions based on the available source code. There are personalization-based strategies (e.g., LG-FedAvg (Liang et al., 2020)), prototype-based approaches (e.g., FedProto (Tan et al., 2022a), FedTGP (Zhang et al., 2024a)), knowledge distillation-based methods (e.g., FML (Shen et al., 2023), FedKD (Wu et al., 2022)), partial-model based model FIARSE (Wu et al., 2024) and method for MHtFGL called FedPPN (Liu et al., 2024b). Due to the need for partial homogeneity of the personalization method, we set up the same classifier for all clients when implementing the LG-FedAvg. In addition, we maintain a completely heterogeneous framework for all client models. The detailed implementations are in Appendix C.
@@ 5.1.3. Heterogeneity Simulations
To partition data to each client, we deploy Louvain (Blondel et al., 2008) community detection algorithm to partition graph into subgraphs that are distributed to clients. To ensure uniformity and simulate multi-clients scenarios, we conduct experiments using configurations of 5, 10, and 20 clients. Specifically, we consider two model heterogeneity scenarios: (1) Divergent model architectures. We utilize five distinct two-layer graph neural network architectures: GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), GIN (Xu et al., 2018a), and SGC (Wu et al., 2019). (2)Varying model scales. We leverage deep GNNs technology  (Xu et al., 2018b) to employ a lightweight two-layer SGC, a conventional two-layer GCN, and deeper models including four-layer GCN, six-layer GCN, and eight-layer GCN. For the allocation of clients, given $K$ clients, the model architecture assigned to the $k$-th client is determined by ($k\bmod 5$), $k\in|K|$. The parameters of each assigned model architecture are re-initialized for every client.
@@ 5.2. Performance Comparison
To answer Q1, we conduct comprehensive experiments under heterogeneous and non-heterogeneous settings. Overall, FedGKC exhibits consistent improvements over all selected MHtFL baselines. (1) Clients with Different Model Architectures. Table 1 summarizes results when clients adopt different model architectures FedGKC achieves the best or second-best performance across all datasets and client configs, surpassing KD and prototype-based methods. Two strong partial-model strategies (i.e., FIARSE, FedPPN) exhibit competitive results due to their effective submodel selection and parameter coverage. However, their performances are less stable when architectural heterogeneity increases, as partial extraction/mapping struggles to preserve topology-aware signals across diverse GNN designs. In contrast, FedGKC decouples cross-client exchange via copilot models and integrates topology-aware distillation, resulting in superior overall accuracy and more robust transfer. (2) Clients with Different Model Scales. In the second scenario, we consider the variations in model scales due to differences in the computational capabilities of various client devices. As shown in Table 2, FedGKC demonstrates superior performance on the test set compared to other methods, with an average of 1.80% higher accuracy than FML and 2.17% higher accuracy than FedKD. It confirms that FedGKC can not only adapt to the heterogeneity of simple model architectures, but can also be extended to models of different scales, and performs well. (3) Comparative Experiments under Non-Heterogeneous Settings. In addition to the experiments under heterogeneous settings, we also conduct comparative experiments in non-heterogeneous settings to evaluate the performance of FedGKC in traditional federated learning environments. The results in Table 3 show that despite being specifically designed for heterogeneous client scenarios, FedGKC performs exceptionally well even in non-heterogeneous settings, outperforming the baseline approach FedAvg by 10.5% on average. FedGKC even surpasses federated learning methods tailored for homogeneous clients by an average of 1.07% over the suboptimal method Fed-Pub, which further validates the generality and robustness of FedGKC.
@@ C.1. Toy Experiment Implementation
The simulation settings and experimental environment is aligned with description in Sec. 5.1. Experiment is conducted on the Cora dataset, which is further partitioned into 10 clients. Representative baselines selected for each category are most competitive MHtFGL methods to our knowledge: FML (Shen et al., 2023) for Partial Model, FedPPN (Liu et al., 2024b) for Prototype, and AlFeCo (Yang et al., 2023) for Public/Synthetic Information.

### v4

@@ 3.4. Knowledge-Aware Model Aggregation
Given the heterogeneity among clients, the learning efficiency of each client’s model varies accordingly. Traditional FL methods typically employ an averaging aggregation strategy (2017FedAvg), which does not take into account the model heterogeneity scenario. This oversight results in conflicted models aggregation and suboptimal overall performance. To address it, we propose a Knowledge-Aware Model Aggregation mechanism that dynamically assigns weights based on learned knowledge quality of each client’s model. Through dynamic weighting, well-informed clients is protected from the adverse effects of clients with limited knowledge. Simultaneously, it channels greater guidance from knowledgeable clients to less-informed ones, improving convergence and robustness.
@@ 5.1.1. Datasets
Experiments are conducted utilizing 8 commonly employed benchmark datasets in graph learning. These consist of three small-scale citation networks (Cora, CiteSeer, PubMed) (CoraDataset), two medium-scale co-authorship networks (CS, Physics) (CSandPhyDataset), two medium-scale user-item datasets (Photo, Computer) (mcauley2015ComputDataset), and one large-scale open graph benchmark dataset (ogbn-arxiv) (ogbnarxiv). Further particulars regarding these datasets are delineated in Table 6.
@@ 5.1.2. Baselines
We compare FedGKC with heterogeneous federation strategies. Given the absence of strategies for MHtFGL, we re-implemented these methods under FGL conditions based on the available source code. There are personalization-based strategies (e.g., LG-FedAvg (liang2020LG)), prototype-based approaches (e.g., FedProto (tan2022fedproto), FedTGP (zhang2024fedtgp)), knowledge distillation-based methods (e.g., FML (shen2023fedfml), FedKD (wu2022communication)), partial-model based model FIARSE (wu2024fiarse) and method for MHtFGL called FedPPN (liu2024model). Due to the need for partial homogeneity of the personalization method, we set up the same classifier for all clients when implementing the LG-FedAvg. In addition, we maintain a completely heterogeneous framework for all client models. The detailed implementations are in Appendix C.
@@ 5.1.3. Heterogeneity Simulations
To partition data to each client, we deploy Louvain (Louvain) community detection algorithm to partition graph into subgraphs that are distributed to clients. To ensure uniformity and simulate multi-clients scenarios, we conduct experiments using configurations of 5, 10, and 20 clients. Specifically, we consider two model heterogeneity scenarios: (1) Divergent model architectures. We utilize five distinct two-layer graph neural network architectures: GCN (kipf2017GCN), GAT (velivckovic2018gat), GraphSAGE (hamilton2017graphsage), GIN (xu2018gin), and SGC (wu2019sgc). (2)Varying model scales. We leverage deep GNNs technology  (jknet) to employ a lightweight two-layer SGC, a conventional two-layer GCN, and deeper models including four-layer GCN, six-layer GCN, and eight-layer GCN. For the allocation of clients, given $K$ clients, the model architecture assigned to the $k$-th client is determined by ($k\bmod 5$), $k\in|K|$. The parameters of each assigned model architecture are re-initialized for every client.
@@ 5.2. Performance Comparison
To answer Q1, we conduct comprehensive experiments under heterogeneous and non-heterogeneous settings. Overall, FedGKC exhibits consistent improvements over all selected MHtFL baselines. (1) Clients with Different Model Architectures. Table 1 summarizes results when clients adopt different model architectures FedGKC achieves the best or second-best performance across all datasets and client configs, surpassing KD and prototype-based methods. Two strong partial-model strategies (i.e., FIARSE, FedPPN) exhibit competitive results due to their effective submodel selection and parameter coverage. However, their performances are less stable when architectural heterogeneity increases, as partial extraction/mapping struggles to preserve topology-aware signals across diverse GNN designs. In contrast, FedGKC decouples cross-client exchange via copilot models and integrates topology-aware distillation, resulting in superior overall accuracy and more robust transfer. (2) Clients with Different Model Scales. In the second scenario, we consider the variations in model scales due to differences in the computational capabilities of various client devices. As shown in Table 2, FedGKC demonstrates superior performance on the test set compared to other methods, with an average of 4.40% higher accuracy than FML and 4.22% higher accuracy than FedKD. It confirms that FedGKC can not only adapt to the heterogeneity of simple model architectures, but can also be extended to models of different scales, and performs well. (3) Comparative Experiments under Non-Heterogeneous Settings. In addition to the experiments under heterogeneous settings, we also conduct comparative experiments in non-heterogeneous settings to evaluate the performance of FedGKC in traditional federated learning environments. The results in Table 3 show that despite being specifically designed for heterogeneous client scenarios, FedGKC performs exceptionally well even in non-heterogeneous settings, outperforming the baseline approach FedAvg by 10.5% on average. FedGKC even surpasses federated learning methods tailored for homogeneous clients by an average of 1.07% over the suboptimal method Fed-Pub, which further validates the generality and robustness of FedGKC.
@@ 6. Conclusion
Outline
To meet page limits, we have included as much essential material as possible in the main text. The appendix provides supplementary details that complete our study. It is organized as follows:
 A:
Algorithm Complexity Analysis of existing FGL studies.
 B:
The Proof of Theorem Analysis presented in Sec. 4.
 C:
Detailed Implementations of Experiment.
Table 5. Algorithm complexity analysis for existing prevalent FL and FGL studies.
  Method Client Memory Server Memory Inference Memory Client Time Server Time Inference Time
FedAvg $\mathcal{O}((b+k)f+f^{2})$ $\mathcal{O}(Nf^{2})$ $\mathcal{O}((b+k)f+f^{2})$ $\mathcal{O}(kmf+nf^{2})$ $\mathcal{O}(N)$ $\mathcal{O}(kmf+nf^{2})$
MOON $\mathcal{O}((b+k)f+Qf^{2})$ $\mathcal{O}(Nf^{2})$ $\mathcal{O}((b+k)f+f^{2})$ $\mathcal{O}(kmf+nf^{2}+Qnf)$ $\mathcal{O}(N)$ $\mathcal{O}(kmf+nf^{2})$
FedSage+ $\mathcal{O}(L((n+sg)f+f^{2}))$ $\mathcal{O}(LtNf^{2})$ $\mathcal{O}(L(n+sg)f+Lf^{2})$ $\mathcal{O}(L((m+sg)f+(n+sg)f^{2}))$ $\mathcal{O}(N)$ $\mathcal{O}(L((m+sg)f+(n+sg)f^{2}))$
FGSSL $\mathcal{O}(Q((b+k)f+f^{2}))$ $\mathcal{O}(Nf^{2})$ $\mathcal{O}(L(b+k)f+Lf^{2})$ $\mathcal{O}(Qkmf+Qnf^{2})$ $\mathcal{O}(N)$ $\mathcal{O}(kmf+nf^{2})$
Fed-PUB $\mathcal{O}(M((b+k)f+f^{2})+M^{2})$ $\mathcal{O}(N(f^{2}+M)+P_{g})$ $\mathcal{O}(M(b+k)f+Mf^{2})$ $\mathcal{O}(Mkmf+Mnf^{2})$ $\mathcal{O}(N^{2}(\log(N)+M^{2}))$ $\mathcal{O}(Mkmf+Mnf^{2})$
FedGTA $\mathcal{O}((b+k)f+f^{2}+kKc)$ $\mathcal{O}(Nf^{2}+NkKc)$ $\mathcal{O}((b+k)f+f^{2})$ $\mathcal{O}(km(f+knc)+n(f^{2}+c))$ $\mathcal{O}(N+NkKc)$ $\mathcal{O}(kmf+nf^{2})$
AdaFGL $\mathcal{O}(E((b+k)f+f^{2}))$ $\mathcal{O}(Nf^{2})$ $\mathcal{O}(E((b+k)f+f^{2}))$ $\mathcal{O}(E(kmf+nf^{2}))$ $\mathcal{O}(N)$ $\mathcal{O}(E(kmf+nf^{2}))$
FedGKC $\mathcal{O}(2(b+l)f+2f^{2})$ $\mathcal{O}(Nf^{2})$ $\mathcal{O}(2(b+l)f+2f^{2})$ $\mathcal{O}(kmf+nf^{2}+f^{2})$ $\mathcal{O}(N)$ $\mathcal{O}(kmf+nf^{2})$
@@ C.1. Toy Experiment Implementation
The simulation settings and experimental environment is aligned with description in Sec. 5.1. Experiment is conducted on the Cora dataset, which is further partitioned into 10 clients. Representative baselines selected for each category are most competitive MHtFGL methods to our knowledge: FML (shen2023fedfml) for Partial Model, FedPPN (liu2024model) for Prototype, and AlFeCo (yang2023allosteric) for Public/Synthetic Information.